# Overcoming the Shadow: Bending Airy Beams for Radiative Near-Field Multi-User Access in Half-Space Blockage Scenarios

### v1

@@ V-E1 Sum-Rate Gain and Interference Suppression
Fig. 5(d) shows that Airy-Opt significantly outperforms both Trad-All (which fails due to blockage) and Airy-Geo (which suffers from interference). A key observation from Fig. 5(f) is the analog coupling strength: the optimized design effectively suppresses the interference leakage $|h_{21}|^{2}$ from the shadowed UE1 to the bright UE2.
@@ V-E2 Physical Mechanism: Service Rebalancing
To understand the root cause of this gain, we analyze the transverse field intensity distribution at the depth of UE2 ($z=300\lambda$), as visualized in Fig. 6. The comparison between Airy-Geo (gray curve) and Airy-Opt (red curve) reveals a crucial energy rebalancing effect:
  • At UE1 (Shadow, $x=-5\lambda$): The Airy-Geo beam delivers very weak energy (approx. 32-32 dB) because its trajectory is not optimally bent for this specific blockage geometry. In contrast, the Airy-Opt beam significantly boosts the signal power to 11-11 dB, a massive gain of +21+21 dB. This confirms that the optimization successfully “pulls” the main lobe energy back into the deep shadow region.
  • At UE2 (Bright, $x=+3.5\lambda$): The Airy-Geo beam creates a strong side-lobe peak (8.6-8.6 dB) at the location of the bright user. The Airy-Opt beam, by adjusting its curvature and angle, reduces this interference to 12.3-12.3 dB. While this is a reduction of only $3.7$ dB, it is sufficient to mitigate interference without sacrificing the primary link.
Core Insight: The essence of the optimization is not merely to place a perfect deep null at UE2, but to rebalance the service between the two users. The geometric solution (Airy-Geo) is overly biased towards the bright region (due to the diffraction nature), starving the shadowed user. The optimized solution (Airy-Opt) sacrifices a small amount of bright-region energy (which is abundant) to critically resurrect the shadowed link, thereby maximizing the overall system throughput.
Figure 6: Comparison of normalized transverse field intensity at the depth of UE2 ($z=300\lambda$). The curves show the beam profiles for the geometric solution (Airy-Geo, gray) and the optimized solution (Airy-Opt, red). The optimization effectively “pulls” energy back into the shadow region to boost UE1 (+21+21 dB) while slightly reducing the interference level at the bright-region user UE2 (3.7-3.7 dB), thereby rebalancing the service quality.
Figure 7: Robustness analysis against lateral positioning errors of the bright-region user (UE2). The beamformer design assumes a nominal position $\hat{\mathbf{p}}_{2}$, while the true position varies by $\Delta x_{2}$. (a) Sum-rate $R_{\text{sum}}$ versus positioning error $\Delta x_{2}/\lambda$: Airy-Opt maintains a robust rate floor across the entire error range. (b) Condition number $\kappa(\mathbf{H}_{\text{eff}})$ versus $\Delta x_{2}/\lambda$: Airy-Opt exhibits a consistently low and flat $\kappa$, indicating stable channel separability. (c) Sum-rate gain relative to the traditional baseline: Airy-Opt provides a sustained gain of >3.5>3.5\,bps/Hz. (d) Worst-case sum-rate (over $\pm\Delta x_{2}$) versus error magnitude: The optimized design significantly outperforms both geometric and traditional benchmarks even under large estimation mismatches.
@@ V-F Robustness to User Positioning Errors
The proposed Airy-null optimization strategy (Section V-C) relies on the spatial distinctness of the users to shape the interference in the analog domain. In practical RNF systems, however, channel state information (CSI) is inevitably imperfect. A critical question arises: Does the fine-tuned Airy beam become hypersensitive to positioning errors, causing the performance to collapse if the user is slightly misaligned? To address this, we evaluate the system’s robustness against lateral positioning errors of the bright-region user (UE2).
@@ V-F1 Error Model and Evaluation Setup
We consider a scenario where the beamformer design is based on an estimated position $\hat{\mathbf{p}}_{2}=(x_{2},z_{2})$, while the actual user is located at $\mathbf{p}_{2}=(x_{2}+\Delta x_{2},z_{2})$. We sweep the lateral error $\Delta x_{2}$ within the range $3\lambda,+3\lambda]$. The setup assumes the “hard case” geometry defined in Section V-A, where the nominal positions are fixed, and the shadowed user (UE1) is assumed to be accurately tracked to isolate the impact of interference leakage.
@@
  • Airy-Geo: UE1 employs the geometric Airy beam (B=25,F=1.75B=-25,F=1.75\,m, $\theta=\theta_{\text{geo}}$) while UE2 uses the traditional beam.
  • Airy-Opt: UE1 employs the optimized Airy beam (B=44,F=1.50B=-44,F=1.50\,m, $\theta=\theta_{\text{geo}}-2.9^{\circ}$) derived from Algorithm 1, while UE2 uses the traditional beam.
Crucially, the precoder for UE2 ($w_{2}$) is fixed to the estimated position $\hat{\mathbf{p}}_{2}$ for all schemes, meaning it does not adapt to the error, thereby creating a realistic mismatch scenario.
@@ V-F2 Results and Physical Interpretation
Gain Retention: Fig. 7(c) illustrates the sum-rate gain of the Airy strategies relative to the traditional baseline. The optimized Airy beam delivers a stable gain (mean +3.89+3.89 bps/Hz) that is largely insensitive to the specific direction of the error. This confirms that the proposed method achieves analog-domain interference shaping, effectively rebalancing the power-interference trade-off in a robust manner. By “pulling” the energy into the shadow (for UE1) while smoothing the side-lobe fluctuations in the bright region (near UE2), the Airy-Opt design ensures that the multi-user cluster remains separable even under realistic estimation uncertainties.

### v2

@@ V-E1 Sum-Rate Gain and Interference Suppression
Fig. 5(d) shows that Airy-Opt significantly outperforms both Trad-All (which fails due to blockage) and Airy-Geo (which suffers from interference). A key observation from Fig. 5(f) is the analog coupling strength: the optimized design effectively suppresses the interference leakage $|h_{21}|^{2}$ from the shadowed UE-1 to the bright UE-2.
@@ V-E2 Physical Mechanism: Service Rebalancing
To understand the root cause of this gain, we analyze the transverse field intensity distribution at the depth of UE-2 ($z=300\lambda$), as visualized in Fig. 6. The comparison between Airy-Geo (gray curve) and Airy-Opt (red curve) reveals a crucial energy rebalancing effect:
  • At UE-1 (Shadow, $x=-5\lambda$): The Airy-Geo beam delivers very weak energy (approx. 32-32 dB) because its trajectory is not optimally bent for this specific blockage geometry. In contrast, the Airy-Opt beam significantly boosts the signal power to 11-11 dB, a massive gain of +21+21 dB. This confirms that the optimization successfully “pulls” the main lobe energy back into the deep shadow region.
  • At UE-2 (Bright, $x=+3.5\lambda$): The Airy-Geo beam creates a strong side-lobe peak (8.6-8.6 dB) at the location of the bright user. The Airy-Opt beam, by adjusting its curvature and angle, reduces this interference to 12.3-12.3 dB. While this is a reduction of only $3.7$ dB, it is sufficient to mitigate interference without sacrificing the primary link.
Core Insight: The essence of the optimization is not merely to place a perfect deep null at UE-2, but to rebalance the service between the two users. The geometric solution (Airy-Geo) is overly biased towards the bright region (due to the diffraction nature), starving the shadowed user. The optimized solution (Airy-Opt) sacrifices a small amount of bright-region energy (which is abundant) to critically resurrect the shadowed link, thereby maximizing the overall system throughput.
Figure 6: Comparison of normalized transverse field intensity at the depth of UE-2 ($z=300\lambda$). The curves show the beam profiles for the geometric solution (Airy-Geo, gray) and the optimized solution (Airy-Opt, red). The optimization effectively “pulls” energy back into the shadow region to boost UE-1 (+21+21 dB) while slightly reducing the interference level at the bright-region user UE-2 (3.7-3.7 dB), thereby rebalancing the service quality.
Figure 7: Robustness analysis against lateral positioning errors of the bright-region user (UE-2). The beamformer design assumes a nominal position $\hat{\mathbf{p}}_{2}$, while the true position varies by $\Delta x_{2}$. (a) Sum-rate $R_{\text{sum}}$ versus positioning error $\Delta x_{2}/\lambda$: Airy-Opt maintains a robust rate floor across the entire error range. (b) Condition number $\kappa(\mathbf{H}_{\text{eff}})$ versus $\Delta x_{2}/\lambda$: Airy-Opt exhibits a consistently low and flat $\kappa$, indicating stable channel separability. (c) Sum-rate gain relative to the traditional baseline: Airy-Opt provides a sustained gain of >3.5>3.5\,bps/Hz. (d) Worst-case sum-rate (over $\pm\Delta x_{2}$) versus error magnitude: The optimized design significantly outperforms both geometric and traditional benchmarks even under large estimation mismatches.
@@ V-F Robustness to User Positioning Errors
The proposed Airy-null optimization strategy (Section V-C) relies on the spatial distinctness of the users to shape the interference in the analog domain. In practical RNF systems, however, channel state information (CSI) is inevitably imperfect. A critical question arises: Does the fine-tuned Airy beam become hypersensitive to positioning errors, causing the performance to collapse if the user is slightly misaligned? To address this, we evaluate the system’s robustness against lateral positioning errors of the bright-region user (UE-2).
@@ V-F1 Error Model and Evaluation Setup
We consider a scenario where the beamformer design is based on an estimated position $\hat{\mathbf{p}}_{2}=(x_{2},z_{2})$, while the actual user is located at $\mathbf{p}_{2}=(x_{2}+\Delta x_{2},z_{2})$. We sweep the lateral error $\Delta x_{2}$ within the range $3\lambda,+3\lambda]$. The setup assumes the “hard case” geometry defined in Section V-A, where the nominal positions are fixed, and the shadowed user (UE-1) is assumed to be accurately tracked to isolate the impact of interference leakage.
@@
  • Airy-Geo: UE-1 employs the geometric Airy beam (B=25,F=1.75B=-25,F=1.75\,m, $\theta=\theta_{\text{geo}}$) while UE-2 uses the traditional beam.
  • Airy-Opt: UE-1 employs the optimized Airy beam (B=44,F=1.50B=-44,F=1.50\,m, $\theta=\theta_{\text{geo}}-2.9^{\circ}$) derived from Algorithm 1, while UE-2 uses the traditional beam.
Crucially, the precoder for UE-2 ($w_{2}$) is fixed to the estimated position $\hat{\mathbf{p}}_{2}$ for all schemes, meaning it does not adapt to the error, thereby creating a realistic mismatch scenario.
@@ V-F2 Results and Physical Interpretation
Gain Retention: Fig. 7(c) illustrates the sum-rate gain of the Airy strategies relative to the traditional baseline. The optimized Airy beam delivers a stable gain (mean +3.89+3.89 bps/Hz) that is largely insensitive to the specific direction of the error. This confirms that the proposed method achieves analog-domain interference shaping, effectively rebalancing the power-interference trade-off in a robust manner. By “pulling” the energy into the shadow (for UE-1) while smoothing the side-lobe fluctuations in the bright region (near UE-2), the Airy-Opt design ensures that the multi-user cluster remains separable even under realistic estimation uncertainties.